# Uncertainty Quantification for Motor Imagery BCI – Machine Learning vs. Deep Learning

### v1

@@ 1 Introduction
Fig. 1: Examples of Event Related Desynchronisation (blue) and Synchronisation (red) trials in a left vs. right hand motor imagery task, with illustrative classifications and probabilities (i.e. confidence) from a model below each. Only when there is a unilateral desychronization over the motor cortex the model is confident and gives a high probability for the correct class, otherwise it is uncertain. Class probabilities should reflect the probability of being correct, and should be useful for separating between correct and incorrect classifications.
The field of Uncertainty Quantification (UQ) focuses on the confidence of ML models when making predictions. Confidence is defined as the predicted probability of a given prediction from an ML to be correct (where ”uncertainty” is simply the opposite of ”confidence”). In practice this is the probabilities you see before applying a threshold to select a class. The goal of UQ is to ensure that the confidence of an ML model aligns with the actual accuracy of the model. If the model has a confidence of 70%, then it should be correct in 70% of the time. To assess this, the calibration error can be measured, which compares the confidence of the ML model to its actual accuracy. Good uncertainty estimation should also allows us to intervene when a given prediction has a low probability of being correct [2]. These desired behaviors are illustrated in Figure 1.
@@ 2 Methods
We our experiments we use a multiple existing dataset using a standard Motor Imagery BCI set-up where participants imagine a cued movement which is then decoded from EEG signals. Specifically, we use the datasets from Steyrl et al. [8], Zhou et al. [9], the BCI Competition IV - dataset 2b [10] and BCI Competition IV - dataset 2a [11]. All datasets were accessed through Mother Of All BCI Benchmarks (MOABB; [12]). We apply a 80-20 within subject train-test split. The Deep Learning models use 10% of the train split as validation data for early stopping. We aim to keep things consistent between methods and datasets by maintaining the same preprocessing pipelines. All data is preprocessed using a single non-causal IIR Band-pass filter, with a frequency band of $7.5-30$Hz. The full code implementation can be found at github.com/Jorissuurmeijer/UQ-motor-imagery.
@@ 2.1 Uncertainty Quantification Methods
Minimum Distance to Riemannian Mean (MDRM; [5]) is a distance-based classifier that relies on geometry of covariance matrices, and is well-established in BCI literature. For MDRM each epoch is represented as the covariance matrix between EEG channels, and we learn a mean covariance matrix for each class. The covariance matrices fall on a manifold, and we use Riemannian Geometry to measure distance over this manifold, instead of through euclidean space. Inference is done by finding the Riemannian distances to each class mean, where the closest class is the prediction. To get class probabilities (i.e. confidence) the softmax function is used.
@@
By scaling down all the logits with large temperature $T$ all the predicted class probabilities come closer together. This means the models predictions will be more uncertain. With a small temperature $T$ the probabilities are pushed more towards the extremes (0 and 1) and thus the models predictions will be more confident. The temperature parameter $T$ is optimized after training the model by finding the value that minimizes the calibration error on the training data. This way the model should be neither overconfident nor underconfident.
Common Spatial Patterns with Linear Discriminant Analysis (CSP-LDA; [4]) is a combination of two Machine Learning methods commonly used in EEG classification. CSP learns a linear combinations of channels (filters) to form surrogate sensors. Each filter (here: 8) maximizes the variance for one class while minimizing the variance for the other class(es). We then take the average band power so that each filter becomes a scalar feature for an LDA classifier. LDA finds an optimal axis that samples can then be projected onto, to achieve class separation. Class probabilities are calculated using Bayes’ rule by assuming each class follows a multivariate Gaussian distribution.
@@
Deep Ensembles [15] uses five ShallowConvNet models trained on the same data. Through differences in random initialization these models will make slightly different predictions, which increases the uncertainty. This allows the model the represent model uncertainty, which is what arises when a model needs to make inference on data that is dissimilar to what it was trained on. Deep Ensembles are often shown to be the best-performing Uncertainty Quantification method in Deep Learning [3].
@@ 2.2.1 Rejection ability
Good uncertainty estimation should be able to distinguish between correct and incorrect predictions. This would allow a model with good UQ to reject samples that are difficult, and only give predictions for samples in which it is confident. To evaluate this we use rejection-accuracy plots. These are created by sorting the test-samples by model uncertainty, and rejecting the most uncertain samples. The accuracy is then evaluated on the remaining samples. By rejecting an increasing number of samples, a model with good UQ should be able to increase its accuracy.
@@ 3 Results
Table 1: Comparison of within-subject mean performance, with the standard deviation across subjects. Boldface indicates the best performing model for a metric and dataset. Deep Learning methods (DE, CNN) give better classifications (Acc. & Brier), but that MDRM-T and CSP-LDA give better uncertainties (NCE & ECE). Deep Learning is more computationally expensive.
@@
Table 1 provides a comparison of all metrics of all models over all datasets. We see that the Deep Ensemble (DE) and the CNN generally give the highest accuracy with negligable difference between them. Only on the dataset from Steyrl [8] the CSP-LDA gives better accuracy. We see that MDRM and MDRM-T have the same accuracies, because the classifications are not affected by Temperature Scaling.
@@
We also consider the difference in computational time between the models, because Deep Learning models are sometimes considered a poor fit for online continual BCI systems because of their increased computational cost. We observed the average training time (per subject) and average inference time (per sample) using a Macbook M1 Pro (2021, 10 core CPU/16 core GPU, 32GB RAM). Deep Learning models can take 30-150 seconds to train while the non-Deep Learning models are trained within one second. Inference time per sample for Deep Learning is over 1 millisecond, while the non-Deep Learning models take less than 0.1 millisecond. This shows that Deep Learning systems are much more computationally expensive, but on a modern high-end laptop using ShallowConvNet [7] the extra training time is not prohibitive, and the extra inference time is not a problem for standard setups.
Lastly we consider the ability of the uncertainty to be able to distinguish between easy and difficult samples. This is done through the rejection-accuracy plots shown in Figure 3. We can see that by rejecting the most uncertain samples all models are able to achieve higher accuracies on all datasets.
@@ 5 Conclusion
This study demonstrates how to systematically analyze in-distribution uncertainty estimation performance in Motor-Imagery BCIs, looking at ECE, NCE, Brier Score and rejection ability. This is a clear and simple setup than can be applied to future models to ensure the predicted class probabilities are reliable. This setup is task-agnostic and evaluates uncertainty in general, so the findings from the experiments that we show here can for a basis for specific uncertainty quantification tasks for Motor Imagery BCIs [1, 18].

### v2

@@ 1 Introduction
Fig. 1: Examples of Event Related Desynchronisation (blue) and Synchronization (red) trials in a left vs. right hand motor imagery task, with illustrative classifications and probabilities (i.e. confidence) from a model below each. Only when there is a unilateral desynchronization over the motor cortex the model is confident and gives a high probability for the correct class, otherwise it is uncertain. Class probabilities should reflect the probability of being correct, and should be useful for separating between correct and incorrect classifications.
The field of Uncertainty Quantification (UQ) focuses on the confidence of ML models when making predictions. Confidence is defined as the predicted probability of a given prediction from an ML to be correct (where ”uncertainty” is simply the opposite of ”confidence”). In practice this is the probability you see before applying a threshold to select a class. The goal of UQ is to ensure that the confidence of an ML model aligns with the actual accuracy of the model. If the model has a confidence of 70%, then it should be correct in 70% of the time. To assess this, the calibration error can be measured, which compares the confidence of the ML model to its actual accuracy. Good uncertainty estimation should also allow us to intervene when a given prediction has a low probability of being correct [2]. These desired behaviors are illustrated in Figure 1.
@@ 2 Methods
In our experiments we use multiple existing datasets from a standard Motor Imagery BCI set-up where participants imagine a cued movement which is then decoded from EEG signals. Specifically, we use the datasets from Steyrl et al. [8], Zhou et al. [9], the BCI Competition IV - dataset 2b [10] and BCI Competition IV - dataset 2a [11]. All datasets were accessed through Mother Of All BCI Benchmarks (MOABB; [12]). We apply an 80-20 within-subject train-test split. The Deep Learning models use 10% of the train split as validation data for early stopping. We aim to keep things consistent between methods and datasets by maintaining the same preprocessing pipelines. All data is preprocessed using a single non-causal IIR Band-pass filter, with a frequency band of $7.5-30$Hz. The full code implementation can be found at github.com/Jorissuurmeijer/UQ-motor-imagery.
@@ 2.1 Uncertainty Quantification Methods
Minimum Distance to Riemannian Mean (MDRM; [5]) is a distance-based classifier that relies on the geometry of covariance matrices, and is well-established in BCI literature. For MDRM each epoch is represented as the covariance matrix between EEG channels, and we learn a mean covariance matrix for each class. The covariance matrices fall on a manifold, and we use Riemannian Geometry to measure distance over this manifold, instead of through euclidean space. Inference is done by finding the Riemannian distances to each class mean, where the closest class is the prediction. To get class probabilities (i.e. confidence) the softmax function is used.
@@
By scaling down all the logits with a large temperature $T$ all the predicted class probabilities come closer together. This means the model’s predictions will be more uncertain. With a small temperature $T$ the probabilities are pushed more towards the extremes (0 and 1) and thus the model’s predictions will be more confident. The temperature parameter $T$ is optimized after training the model by finding the value that minimizes the calibration error on the training data. This way the model should be neither overconfident nor underconfident.
Common Spatial Patterns with Linear Discriminant Analysis (CSP-LDA; [4]) is a combination of two Machine Learning methods commonly used in EEG classification. CSP learns linear combinations of channels (filters) to form surrogate sensors. Each filter (here: 8) maximizes the variance for one class while minimizing the variance for the other class(es). We then take the average band power so that each filter becomes a scalar feature for an LDA classifier. LDA finds an optimal axis that samples can then be projected onto, to achieve class separation. Class probabilities are calculated using Bayes’ rule by assuming each class follows a multivariate Gaussian distribution.
@@
Deep Ensembles [15] uses five ShallowConvNet models trained on the same data. Through differences in random initialization these models will make slightly different predictions, which increases the uncertainty. This allows the model to represent model uncertainty, which is what arises when a model needs to make inferences on data that is dissimilar to what it was trained on. Deep Ensembles are often shown to be the best-performing Uncertainty Quantification method in Deep Learning [3].
@@ 2.2.1 Rejection ability
Good uncertainty estimation should be able to distinguish between correct and incorrect predictions. This would allow a model with good UQ to reject samples that are difficult, and only give predictions for samples in which it is confident. To evaluate this we use rejection-accuracy plots. These are created by sorting the test samples by model uncertainty, and rejecting the most uncertain samples. The accuracy is then evaluated on the remaining samples. By rejecting an increasing number of samples, a model with good UQ should be able to increase its accuracy.
@@ 3 Results
Table 1: Comparison of within-subject mean performance, with the standard deviation across subjects. Boldface indicates the best-performing model for a metric and dataset. Deep Learning methods (DE, CNN) give better classifications (Acc. & Brier), but MDRM-T and CSP-LDA give better uncertainties (NCE & ECE). Deep Learning is more computationally expensive.
@@
Table 1 provides a comparison of all metrics of all models over all datasets. We see that the Deep Ensemble (DE) and the CNN generally give the highest accuracy with negligible difference between them. Only on the dataset from Steyrl [8] the CSP-LDA gives better accuracy. We see that MDRM and MDRM-T have the same accuracies, because the classifications are not affected by Temperature Scaling.
@@
We also consider the difference in computational time between the models, because Deep Learning models are sometimes considered a poor fit for online continual BCI systems because of their increased computational cost. We observed the average training time (per subject) and average inference time (per sample) using a MacBook M1 Pro (2021, 10-core CPU/16-core GPU, 32GB RAM). Deep Learning models can take 30-150 seconds to train while the non-Deep Learning models are trained within one second. Inference time per sample for Deep Learning is over 1 millisecond, while the non-Deep Learning models take less than 0.1 millisecond. This shows that Deep Learning systems are much more computationally expensive, but on a modern high-end laptop using ShallowConvNet [7] the extra training time is not prohibitive, and the extra inference time is not a problem for standard setups.
Lastly we consider the ability of the uncertainty to distinguish between easy and difficult samples. This is done through the rejection-accuracy plots shown in Figure 3. We can see that by rejecting the most uncertain samples all models are able to achieve higher accuracies on all datasets.
@@ 5 Conclusion
This study demonstrates how to systematically analyze in-distribution uncertainty estimation performance in Motor-Imagery BCIs, looking at ECE, NCE, Brier Score and rejection ability. This is a clear and simple setup that can be applied to future models to ensure the predicted class probabilities are reliable. This setup is task-agnostic and evaluates uncertainty in general, so the findings from the experiments that we show here can form a basis for specific uncertainty quantification tasks for Motor Imagery BCIs [1, 18].